# Ring formation from black hole superradiance through repeated particle production on bound orbits

### v1

@@ III.1 The classical dynamics and the formation of the particle ring
The $\chi$ particles are produced as localized wave packets from the resonance region of the $\phi$ cloud. Since their de Broglie wavelength is negligible compared to the size of the cloud, their subsequent evolution can be analyzed by treating them as classical, point-like particles. To this end, we employ the WKB approximation, substituting the ansatz $\chi=A(\mathbf{r},t)e^{iS(\mathbf{r},t)}$ into the Klein-Gordon equation:
$[\nabla^{\mu}\nabla_{\mu}-(\mu_{\chi}^{2}+\lambda\phi)]\chi=0\,.$ (15)
@@
where $\gamma=\sqrt{1-\dot{\mathbf{r}}^{2}}$ is the Lorentz factor. The resulting equation of motion is the key to understanding the post-production fate of the $\chi$ particles. It describes the particle’s trajectory under the combined influence of the weak gravitational potential $\Phi_{N}$ and the non-gravitational force from the gradient of the $\phi$ cloud, $\sim-\nabla m_{\mathrm{eff}}$.
@@
Figure 1: The radial motion of a particle in a numerically solved, stable quasi-circular orbit, plotted over one orbital period with benchmark parameters ($\mu_{\chi}/\mu_{\phi}=100$, $\lambda\Phi_{0}/\mu_{\chi}^{2}=1$, and $\alpha=0.15$). Initial conditions are $r_{i}=2r_{0}$, $\theta_{i}=\pi/2$, and $\varphi_{i}=\pi+0.1$, modeling the particle initially produced near the resonance region with angular velocity $\dot{\varphi}_{i}\simeq 1.24\alpha$ and initial momenta $k\simeq 0.2\mu_{\chi}$. The radial coordinate $r$ oscillates with only small deviations around the initial radius $2r_{0}$, confirming the stability of the orbit.
@@
Figure 2: The illustration of the sustained amplification mechanism for a trapped $\chi$ particle. This figure schematically shows the “staircase” growth of a trapped mode’s occupation number |$\absolutevalue{n_{{k}_{\star}}}$ | as it moves along its quasi-circular orbit. The smooth, horizontal segment (black line) represents the particle’s motion between resonant regions, where its amplitude remains constant. The sharp vertical jumps in the blue line represent brief, intense amplification of the occupation number $n_{{k}_{\star}}$ as the particle crosses resonance planes (the gray planes), occurring each time $\phi$ oscillates to its minimum (the phase of $\phi$ is shown by the red line).
@@
For most particles, the initial “kick” from resonance places them on trajectories that are quickly destabilized by the perturbing force, causing them to either escape to infinity or be accreted by the black hole. A distinct dynamical pathway, however, emerges for only a select fraction of $k$ modes generated with initial conditions that favor a quasi-circular orbit. The key lies in the spatially dependent profile of the superradiant cloud, which peaks at the toroidal region. This naturally singles out a special class of particles: those generated with low tangential momentum $k_{\star}\sim\alpha\mu_{\chi}$. This condition makes sure that the particles periodically pass through the resonance region and generate more particles on the same orbit. Over time, these selected modes become dominant, as other particles—produced initially—follow trajectories that quickly take them out of the optimal resonance region.
@@
While the non-gravitational force from the cloud, $\sim-\nabla m_{\mathrm{eff}}$, acts as a force that would normally destabilize a simple Keplerian orbit, a stable quasi-circular path is nevertheless possible. This is due to a crucial separation of timescales: the force from the $\phi$ field oscillates rapidly at frequency $\omega\sim\mu_{\phi}$, whereas the particle’s orbital frequency is much lower, i.e., $\Omega\sim\alpha^{2}\omega\ll\omega$. Over a single orbit, the effects of this rapidly oscillating force are largely averaged out, allowing the particle to maintain its near-circular trajectory. Our numerical solutions to Eq. (19) confirm that such trajectories are indeed possible, as illustrated in Fig. 1, which shows a particle’s radial coordinate remaining stable with slight oscillations. The long-term stability of these orbits is further confirmed by a formal perturbative analysis detailed in Appendix C.
@@
This orbital configuration enables the mechanism of sustained amplification. Particles that enter the quasi-circular orbit will repeatedly traverse the resonant regions. The number of crossings per orbit is substantial, i.e., $N_{\text{crossings}}={\mu_{\phi}}/{\Omega}\sim{\alpha^{-2}}$. This implies that approximately $\alpha^{-2}$ particle production events occur per orbital cycle, demonstrating that amplification efficiency in such orbits far exceeds all other orbital configurations.
@@
This hybrid process—classical orbital motion punctuated by bursts of amplification—is what leads to the “staircase” growth of the particle’s occupation number, as schematically depicted in Fig. 2. These frequent, periodic amplifications provide a unique mechanism for particles in these select orbits to continuously extract energy from the $\phi$ cloud. In contrast, other particles on different trajectories rapidly diminish in the fraction of the total $\chi$ particles. This sustained energy input drives the virialization of the trapped particles, causing them to settle into a quasi-steady-state configuration. The final structure, as hypothesized, is a toroidal ring located in the region of most efficient production and amplification: a circular band around $r\approx 2r_{0}$ in the equatorial plane. Note that our WKB-based description of a classical particle on a well-defined trajectory is valid during the orbital periods between resonance events; in this phase, particle number is conserved. The classical picture breaks down only for the brief moments when the particle traverses the resonance planes. At these instances, the process is properly described as an amplification of the $\chi$ field. Since the duration of each amplification event is negligible compared to the overall orbital period, these periodic “kicks” do not significantly alter the long-term trajectory. Therefore, the particle’s overall path remains well-described by the effective classical equation of motion, Eq. (19).
@@ III.2 Backreaction and Saturation
The sustained amplification described in the previous section provides a powerful mechanism for transferring energy from the $\phi$ cloud to the $\chi$ ring. This process, however, cannot continue indefinitely. As the $\chi$ ring grows in mass and density, its backreaction on the primary $\phi$ cloud becomes significant, eventually halting the energy transfer. In this section, we analyze this backreaction process to determine the final, saturated state of the system.
Table 1: Benchmark parameters for the saturated cloud-ring system. We present four representative cases: two for a stellar-mass black hole ($M_{\text{BH}}=12.5M_{\odot}$) and two for a supermassive black hole ($M_{\text{BH}}=2.5\times 10^{9}M_{\odot}$). For each, we show results for an early saturation (smaller $\Phi_{0c}$) versus a late saturation (larger $\Phi_{0c}$) scenario. The specific parameters used for these scenarios are as follows. For the stellar-mass case, we use $\alpha=0.1$, $\mu_{\phi}=10^{-12}\,\mathrm{eV}$, $\mu_{\chi}=10^{-10}\,\mathrm{eV}$ with $\lambda=10^{-40}\,\mathrm{eV}$ (for the early saturation case) and $\lambda=10^{-45}\,\mathrm{eV}$ (for the late saturation case). For the SMBH case, we use $\alpha=0.2$, $\mu_{\phi}=10^{-20}\,\mathrm{eV}$, $\mu_{\chi}=10^{-17}\,\mathrm{eV}$ with $\lambda=10^{-54}\,\mathrm{eV}$ (for the early saturation case) and $\lambda=10^{-59}\,\mathrm{eV}$ (for the late saturation case).
@@
Stellar BH (Late Sat.) $12.5M_{\odot}$ $10^{16}\text{GeV}$ $16.67M_{\odot}$ $0.001M_{\odot}$ $10^{73}$ $6\times 10^{-5}$
@@
Since the source term $\left\langle\chi^{2}\right\rangle$ contains rapidly oscillating terms and their frequencies are much higher than $\mu_{\phi}$, we can safely apply the oscillation averaged expectation value $\frac{1}{2}\lambda\expectationvalue{\chi^{2}}$ to estimate the source term in Eq. (24). After this averaging, the term can be well-approximated by the local number density (see Appendix D for a detailed derivation):
@@
The backreaction becomes significant when this new source term becomes comparable to the original mass term in the equation, i.e., $\frac{1}{2}\lambda\expectationvalue{\chi^{2}}\sim\mu_{\phi}^{2}\phi$, as treated in Ref. Kofman et al. (1997). From this condition, we can estimate the critical number density, $n_{\chi,c}$, at which the growth of the ring saturates:
@@
Building on this, we can now estimate the total mass of the saturated $\chi$ ring, $M_{\chi}$. The total mass of the ring is found by integrating the critical number density over the volume ($M_{\chi}=\mu_{\chi}n_{\chi,c}\mathcal{V}$). We approximate the ring as a torus, estimating its volume as the product of its major circumference $2\pi r_{0}$ and its cross-sectional area taken as $\pi k_{\star}^{-1}$ with $k_{\star}\sim\alpha\mu_{\chi}$ being the characteristic momentum of the trapped particles. A detailed calculation yields:
@@ IV Conclusion and Discussion
In this work, we have explored a novel phenomenology that emerges from a coupled axion-like particles system undergoing superradiance around a Kerr black hole. We demonstrated that the coupling term, $\frac{1}{2}\lambda\phi\chi^{2}$, can trigger the repeated resonant production of the heavier particles, $\chi$, from the primary superradiant cloud of a lighter scalar, $\phi$. By analyzing the full dynamical sequence—from the initial growth of the $\phi$ cloud to the subsequent particle production—we identified a novel mechanism for the formation of a stable structure of $\chi$. We showed that particles produced with specific low, tangential momenta can become gravitationally trapped. Using a WKB analysis, we found that these particles follow quasi-stable circular orbits, allowing them to be repeatedly amplified by the oscillating $\phi$ cloud. This sustained energy injection drives the system towards a saturated equilibrium, resulting in the formation of a stable ring of $\chi$ particles embedded within the primary cloud. Distinct from other studies of resonant particle production, this work takes into consideration the spatiotemporal onset of the resonance. We have shown that particle production does not occur uniformly throughout the cloud. Instead, it commences at the location where $\phi$ field’s amplitude first reaches the resonant threshold—the toroidal peak of the wavefunction at $r=2r_{0}$. It is this localized ignition of particle production that naturally singles out the circular orbit at this specific radius, which allows the $\chi$ particles to cross the resonant region repetitively, providing a robust physical reason for the ring’s formation site and its subsequent stability. This crucial link between the cloud’s spatial profile and the resulting particle dynamics has not been emphasized in previous literature.
@@
Furthermore, we briefly address the contrasting case of the quartic interaction, $\frac{1}{2}g^{2}\phi^{2}\chi^{2}$. While this coupling also triggers parametric resonance, it leads to a more violent and dissipative outcome. The extreme conditions require the interaction energy term $g^{2}\Phi_{0}^{2}$ to far exceed the particle’s bare mass squared, $\mu_{\chi}^{2}$. Consequently, the post-production dynamics are governed by a powerful non-gravitational force proportional to $g^{2}\nabla^{2}\phi^{2}$, which completely dominates the particle’s trajectory. Instead of settling into stable orbits, the produced particles are rapidly accelerated to relativistic energies and ejected from the system. This scenario is analogous to the fermion acceleration by the scalar cloud discussed in Ref. Chen et al. (2025). The final state is therefore not a stable ring but a powerful, escaping flux that acts as a dissipative channel, quenching the cloud’s growth. A detailed analysis of the resonance condition and particle acceleration for this case is deferred to Appendix A.
@@
The final, stable cloud-ring system may also offer new phenomenological possibilities. In the presence of a binary companion, for instance, it is known that resonance transition from tidal perturbations can deplete the primary $\phi$ cloud efficiently Baumann et al. (2019b); Berti et al. (2019). Our work suggests a potential consequence of such an event: the rapid depletion of the $\phi$ cloud could alter the dynamical environment for the secondary $\chi$ ring, possibly causing it to evolve into a more diffuse, extended halo over time. The presence of this new halo of heavy particles might then leave a distinct imprint on the subsequent gravitational wave signal from the binary inspiral, providing an indirect but powerful probe of this phenomenology.

### v2

@@ III.1 The classical dynamics and the formation of the particle ring
The $\chi$ particles are produced as localized wave packets from the resonance region of the $\phi$ cloud. Since their de Broglie wavelength is negligible compared to the size of the cloud, their subsequent evolution can be analyzed by treating them as classical, pointlike particles. To this end, we employ the WKB approximation, substituting the ansatz $\chi=A(\mathbf{r},t)e^{iS(\mathbf{r},t)}$ into the Klein-Gordon equation:
$\nabla^{\mu}\nabla_{\mu}-(\mu_{\chi}^{2}+\lambda\phi)]\chi=0\,.$ (15)
@@
where $\gamma=\sqrt{1-\dot{\mathbf{r}}^{2}}$ is the Lorentz factor. The resulting equation of motion is the key to understanding the postproduction fate of the $\chi$ particles. It describes the particle’s trajectory under the combined influence of the weak gravitational potential $\Phi_{N}$ and the nongravitational force from the gradient of the $\phi$ cloud, $\sim-\nabla m_{\mathrm{eff}}$.
@@
Figure 1: The radial motion of a particle in a numerically solved, stable quasicircular orbit, plotted over one orbital period with benchmark parameters ($\mu_{\chi}/\mu_{\phi}=100$, $\lambda\Phi_{0}/\mu_{\chi}^{2}=1$, and $\alpha=0.15$). Initial conditions are $r_{i}=2r_{0}$, $\theta_{i}=\pi/2$, and $\varphi_{i}=\pi+0.1$, modeling the particle initially produced near the resonance region with angular velocity $\dot{\varphi}_{i}\simeq 1.24\alpha$ and initial momenta $k\simeq 0.2\mu_{\chi}$. The radial coordinate $r$ oscillates with only small deviations around the initial radius $2r_{0}$, confirming the stability of the orbit.
@@
Figure 2: The illustration of the cumulative amplification mechanism for a trapped $\chi$ particle. This figure schematically shows the “staircase” growth of a trapped mode’s occupation number $|\absolutevalue{n_{{k}_{\star}}}$ as it moves along its quasicircular orbit. The smooth, horizontal segment (black line) represents the particle’s motion between resonant regions, where its amplitude remains constant. The sharp vertical jumps in the blue line represent brief, intense amplification of the occupation number $n_{{k}_{\star}}$ as the particle crosses resonance planes (the gray planes), occurring each time $\phi$ oscillates to its minimum (the phase of $\phi$ is shown by the red line).
@@
For most particles, the initial “kick” from resonance places them on trajectories that are quickly destabilized by the perturbing force, causing them to either escape to infinity or be accreted by the black hole. A distinct dynamical pathway, however, emerges for only a select fraction of $k$ modes generated with initial conditions that favor a quasicircular orbit. The key lies in the spatially dependent profile of the superradiant cloud, which peaks at the toroidal region. This naturally singles out a special class of particles: those generated with low tangential momentum $k_{\star}\sim\alpha\mu_{\chi}$. This condition makes sure that the particles periodically pass through the resonance region and generate more particles on the same orbit. Over time, these selected modes become dominant, as other particles—produced initially—follow trajectories that quickly take them out of the optimal resonance region.
@@
While the nongravitational force from the cloud, $\sim-\nabla m_{\mathrm{eff}}$, acts as a force that would normally destabilize a simple Keplerian orbit, a stable quasicircular path is nevertheless possible. This is due to a crucial separation of timescales: the force from the $\phi$ field oscillates rapidly at frequency $\omega\sim\mu_{\phi}$, whereas the particle’s orbital frequency is much lower, i.e., $\Omega\sim\alpha^{2}\omega\ll\omega$. Over a single orbit, the effects of this rapidly oscillating force are largely averaged out, allowing the particle to maintain its near-circular trajectory. Our numerical solutions to Eq. (19) confirm that such trajectories are indeed possible, as illustrated in Fig. 1, which shows a particle’s radial coordinate remaining stable with slight oscillations. The long-term stability of these orbits is further confirmed by a formal perturbative analysis detailed in Appendix C.
@@
This orbital configuration enables the mechanism of cumulative amplification. Particles that enter the quasicircular orbit will repeatedly traverse the resonant regions. The number of crossings per orbit is substantial, i.e., $N_{\text{crossings}}={\mu_{\phi}}/{\Omega}\sim{\alpha^{-2}}$. This implies that approximately $\alpha^{-2}$ particle production events occur per orbital cycle, demonstrating that amplification efficiency in such orbits far exceeds all other orbital configurations.
@@
This hybrid process—classical orbital motion punctuated by bursts of amplification—is what leads to the “staircase” growth of the particle’s occupation number, as schematically depicted in Fig. 2. These frequent, periodic amplifications provide a unique mechanism for particles in these select orbits to continuously extract energy from the $\phi$ cloud. In contrast, other particles on different trajectories rapidly diminish in the fraction of the total $\chi$ particles. This cumulative energy input drives the virialization of the trapped particles, causing them to settle into a quasi-steady-state configuration. The final structure, as hypothesized, is a toroidal ring located in the region of most efficient production and amplification: a circular band around $r\approx 2r_{0}$ in the equatorial plane. Note that our WKB-based description of a classical particle on a well-defined trajectory is valid during the orbital periods between resonance events; in this phase, particle number is conserved. The classical picture breaks down only for the brief moments when the particle traverses the resonance planes. At these instances, the process is properly described as an amplification of the $\chi$ field. Since the duration of each amplification event is negligible compared to the overall orbital period, these periodic “kicks” do not significantly alter the long-term trajectory. Therefore, the particle’s overall path remains well described by the effective classical equation of motion, Eq. (19).
@@ III.2 Backreaction and saturation
The cumulative amplification described in the previous section provides a powerful mechanism for transferring energy from the $\phi$ cloud to the $\chi$ ring. This process, however, cannot continue indefinitely. As the $\chi$ ring grows in mass and density, its backreaction on the primary $\phi$ cloud becomes significant, eventually halting the energy transfer. In this section, we analyze this backreaction process to determine the final, saturated state of the system.
Table 1: Benchmark parameters for the saturated cloud-ring system. We present four representative cases: two for a stellar-mass black hole ($M_{\text{BH}}=12.5M_{\odot}$) and two for a supermassive black hole ($M_{\text{BH}}=2.5\times 10^{9}M_{\odot}$). For each, we show results for an early saturation (smaller $\Phi_{0c}$) versus a late saturation (larger $\Phi_{0c}$) scenario. The specific parameters used for these scenarios are as follows. For the stellar-mass case, we use $\alpha=0.1$, $\mu_{\phi}=10^{-12}\,\mathrm{eV}$, $\mu_{\chi}=10^{-10}\,\mathrm{eV}$ with $\lambda=10^{-40}\,\mathrm{eV}$ (for the early saturation case) and $\lambda=10^{-44}\,\mathrm{eV}$ (for the late saturation case). For the SMBH case, we use $\alpha=0.2$, $\mu_{\phi}=10^{-20}\,\mathrm{eV}$, $\mu_{\chi}=10^{-17}\,\mathrm{eV}$ with $\lambda=10^{-54}\,\mathrm{eV}$ (for the early saturation case) and $\lambda=10^{-59}\,\mathrm{eV}$ (for the late saturation case).
@@
Stellar BH (Late Sat.) $12.5M_{\odot}$ $10^{15}\text{GeV}$ $0.167M_{\odot}$ $10^{-5}M_{\odot}$ $10^{71}$ $6\times 10^{-5}$
@@
Since the source term $\left\langle\chi^{2}\right\rangle$ contains rapidly oscillating terms and their frequencies are much higher than $\mu_{\phi}$, we can safely apply the oscillation averaged expectation value $\frac{1}{2}\lambda\expectationvalue{\chi^{2}}$ to estimate the source term in Eq. (24). After this averaging, the term can be well approximated by the local number density (see Appendix D for a detailed derivation):
@@
The backreaction becomes significant when this new source term becomes comparable to the original mass term in the equation, i.e., $\frac{1}{2}\lambda\expectationvalue{\chi^{2}}\sim\mu_{\phi}^{2}\phi$, as treated in Ref. Kofman:1997yn. From this condition, we can estimate the critical number density, $n_{\chi,c}$, at which the growth of the ring saturates,
@@
Building on this, we can now estimate the total mass of the saturated $\chi$ ring, $M_{\chi}$. The total mass of the ring is found by integrating the critical number density over the volume ($M_{\chi}=\mu_{\chi}n_{\chi,c}\mathcal{V}$). We approximate the ring as a torus, estimating its volume as the product of its major circumference $2\pi r_{0}$ and its cross-sectional area taken as $\pi k_{\star}^{-1}$ with $k_{\star}\sim\alpha\mu_{\chi}$ being the characteristic momentum of the trapped particles. A detailed calculation yields
@@ IV Conclusion and Discussion
In this work, we have explored a novel phenomenology that emerges from a coupled axionlike particles system undergoing superradiance around a Kerr black hole. We demonstrated that the coupling term, $\frac{1}{2}\lambda\phi\chi^{2}$, can trigger the repeated resonant production of the heavier particles, $\chi$, from the primary superradiant cloud of a lighter scalar, $\phi$. By analyzing the full dynamical sequence—from the initial growth of the $\phi$ cloud to the subsequent particle production—we identified a novel mechanism for the formation of a stable structure of $\chi$. We showed that particles produced with specific low, tangential momenta can become gravitationally trapped. Using a WKB analysis, we found that these particles follow quasistable circular orbits, allowing them to be repeatedly amplified by the oscillating $\phi$ cloud. This cumulative energy injection drives the system towards a saturated equilibrium, resulting in the formation of a stable ring of $\chi$ particles embedded within the primary cloud. Distinct from other studies of resonant particle production, this work takes into consideration the spatiotemporal onset of the resonance. We have shown that particle production does not occur uniformly throughout the cloud. Instead, it commences at the location where $\phi$ field’s amplitude first reaches the resonant threshold—the toroidal peak of the wave function at $r=2r_{0}$. It is this localized ignition of particle production that naturally singles out the circular orbit at this specific radius, which allows the $\chi$ particles to cross the resonant region repetitively, providing a robust physical reason for the ring’s formation site and its subsequent stability. This crucial link between the cloud’s spatial profile and the resulting particle dynamics has not been emphasized in previous literature.
A natural question is whether particle-particle scattering processes could disrupt this stable ring structure after its formation. We have analyzed the most relevant channels and found their effects to be negligible. The annihilation of two heavy $\chi$ particles into a much lighter $\phi$ quanta is kinematically forbidden due to the large mass hierarchy ($\mu_{\chi}\gg\mu_{\phi}$), which in turn suppresses any subsequent scattering involving $\phi$ quanta. Furthermore, while elastic scattering $\chi\chi\to\chi\chi$ is possible, its rate is extremely low. The mean free path for these collisions scales with the coupling as $L\propto\lambda^{-2}$ and is found to be orders of magnitude larger than the ring’s radius itself. Consequently, scattering events are far too infrequent to collectively disrupt the ring’s stability.
@@
Furthermore, we briefly address the contrasting case of the quartic interaction, $\frac{1}{2}g^{2}\phi^{2}\chi^{2}$. While this coupling also triggers parametric resonance, it leads to a more violent and dissipative outcome. The extreme conditions require the interaction energy term $g^{2}\Phi_{0}^{2}$ to far exceed the particle’s bare mass squared, $\mu_{\chi}^{2}$. Consequently, the postproduction dynamics are governed by a powerful nongravitational force proportional to $g^{2}\nabla^{2}\phi^{2}$, which completely dominates the particle’s trajectory. Instead of settling into stable orbits, the produced particles are rapidly accelerated to relativistic energies and ejected from the system. This scenario is analogous to the fermion acceleration by the scalar cloud discussed in Ref. Chen:2023vkq. The final state is therefore not a stable ring but a powerful, escaping flux that acts as a dissipative channel, quenching the cloud’s growth. A detailed analysis of the resonance condition and particle acceleration for this case is deferred to Appendix A.
@@
The final, stable cloud-ring system may also offer new phenomenological possibilities. In the presence of a binary companion, for instance, it is known that resonance transition from tidal perturbations can deplete the primary $\phi$ cloud efficiently Baumann:2018vus; Berti:2019wnn. Our work suggests a potential consequence of such an event: the rapid depletion of the $\phi$ cloud could alter the dynamical environment for the secondary $\chi$ ring, possibly causing it to evolve into a more diffuse, extended halo over time. The presence of this new halo of heavy particles might then leave a distinct imprint on the subsequent gravitational wave signal from the binary inspiral, providing an indirect but powerful probe of this phenomenology.